# How Well Can Preference Optimization Generalize Under Noisy Feedback?

### v2

@@ 3.2 Analyzing Generalization via Boundary Dynamics
A key focus of our paper is to provide a tractable analysis of GPO’s generalization behavior under practical considerations. Our analytical framework is designed with practicality in mind. Besides taking realistically noisy feedback into account, we consider the generalization of models after finite gradient steps when the loss is within a constant factor of its initial value, extending the approach in Im & Li (2025) to general objective functions and noisy feedback. This scenario closely matches real-world practices, where large language models are often fine-tuned for a finite number of steps to avoid overfitting. For this reason, our analytical approach is different from classical generalization theory, which typically considers overparameterized models that achieve near-optimal loss (Allen-Zhu et al., 2019; Arora et al., 2019; Cao & Gu, 2020; Subramanian et al., 2022) or are independent of the training process (Arora et al., 2018; Lotfi et al., 2022; 2023).
@@ 3.3 Generalization Guarantee
We now characterize the preference distribution in order to provide a tractable analysis and bound the generalization error. Importantly, the features we model are designed to reflect the characteristics of the real-world transformer backbone, ensuring that our theoretical analysis remains grounded in the specific inductive biases and structures that are typical of such models. Specifically, we find as seen in Table 1 that the embeddings after the RMSNorm layer in practical models such as Llama (Zhang & Sennrich, 2019; Touvron et al., 2023; Grattafiori et al., 2024) have near-uniform norm, suggesting that the embedding distribution can be closely modeled by a hyperspherical distribution. In particular, we consider the von Mises-Fisher (vMF) distribution, a classical and important distribution in directional statistics (Mardia & Jupp, 2009), which is analogous to spherical Gaussian distributions for features with unit norms (see Figure 3). The density function of the vMF distribution is given by $\rho(x;\mu,\kappa)=C_{d}(\kappa)e^{\kappa\mu^{\top}x},$ where $\mu$ represents the mean direction and $\kappa$ is the concentration parameter, and $C_{d}(\kappa)$ normalization constant dependent on the dimension $d$ and $\kappa$. We denote the distribution with mean direction $\mu$ and concentration parameter $\kappa$ as $\text{vMF}(\mu,\kappa)$. We also define a normalized concentration parameter $\gamma=\frac{2\kappa}{d}$.
@@
Theorem 3.3 provides practical guidance for deploying preference optimization in real-world systems where noisy human feedback is common. By quantifying how the generalization performance degrades with increasing noise and characterizing the threshold at which learning fails, our results offer practical guidance on the data quantity and quality required for successful preference optimization. In particular, our theoretical insight suggests a practical diagnostic: before applying a preference optimization algorithm, practitioners can first examine the geometric properties of their dataset. If the data exhibits low concentration or poor separation, the theory implies that standard optimization may be highly sensitive to noise, and a noise-aware or robust preference learning algorithm should be preferred. In contrast, if the data is well-structured, even noisy labels may still yield strong generalization with standard methods. Thus, our framework provides a principled way to inform algorithmic choices based on dataset properties.
@@ 4.2 Verification on Real-World Dataset
To further verify our theory on a real-world dataset, we consider the Anthropic Evaluations datasets (Perez et al., 2022). Unlike the HH-RLHF dataset, which already contains approximately 25-30% inherent noise (Wang et al., 2024), the Anthropic Evaluations datasets provide a cleaner baseline at $\epsilon=0$, allowing us to systematically introduce and analyze the full trend of increasing noise levels. We plot the test accuracies of models trained on two behaviors with varying levels of separation across different noise rates and for different objective functions such as DPO, IPO, and SLiC. The level of separation is determined by the difference between the positive mean embedding and the negative mean embedding for each behavior. We train with noise rates ranging from $\epsilon=0$ to $\epsilon=0.5$ with 0.05 increments for mislabeled datasets and corresponding values of $\omega$ for uncertain datasets, and measure the test performance for each setting. We use 90% of the dataset for training and 10% for testing. We perform DPO/IPO/SLiC for 1 epoch on the noisy datasets. We provide the complete training hyperparameters in Appendix A and additional details on uncertain datasets in Appendix D.
@@ 5 Related Works
A key aspect of training and deploying large language models is ensuring the models behave in safe and helpful ways (Ji et al., 2023; Casper et al., 2023; Hendrycks et al., 2021; Leike et al., 2018). This is an important problem due to the potential harms that can arise in large models (Park et al., 2023; Carroll et al., 2023; Perez et al., 2022; Sharma et al., 2023; Bang et al., 2023; Hubinger et al., 2019; Berglund et al., 2023; Ngo et al., 2022; Shevlane et al., 2023; Shah et al., 2022; Pan et al., 2022). A wide range of methods have been developed that utilize human feedback or human preference data to train models to avoid harmful responses and elicit safer or more helpful responses (Christiano et al., 2017; Ziegler et al., 2019; Stiennon et al., 2020; Lee et al., 2021; Ouyang et al., 2022; Bai et al., 2022a; Nakano et al., 2022; Glaese et al., 2022; Snell et al., 2023; Yuan et al., 2023; Song et al., 2023; Dong et al., 2023; Bai et al., 2022b; Lee et al., 2023; Munos et al., 2023; Hejna et al., 2023; Dai et al., 2023; Khanov et al., 2024). Particularly, the Reinforcement Learning from Human Feedback (RLHF) framework has proven effective in aligning large pre-trained language models (Christiano et al., 2017; Ziegler et al., 2019; Ouyang et al., 2022; Bai et al., 2022a). However, given its computational inefficiency, recent shifts in focus favor closed-form losses that directly utilize offline preferences, like Direct Preference Optimization (DPO) (Rafailov et al., 2023) and related methodologies (Azar et al., 2023; Pal et al., 2024; Liu et al., 2024b; Ethayarajh et al., 2024a; Xiong et al., 2023; Tang et al., 2024; Meng et al., 2024; Ethayarajh et al., 2024b; Zeng et al., 2024; Calandriello et al., 2024; Muldrew et al., 2024; Ray Chowdhury et al., 2024; Liu et al., 2024a; Gao et al., 2024a; Yang et al., 2024; Chakraborty et al., 2024). Despite the empirical success and wide adoption in real-world systems (OpenAI, 2023; Anthropic, 2023; Touvron et al., 2023), fewer works provide theoretical underpinnings (Azar et al., 2023; Rafailov et al., 2024; Im & Li, 2024; Tang et al., 2024; Ray Chowdhury et al., 2024; Tajwar et al., 2024; Xu et al., 2024; Nika et al., 2024; Xiong et al., 2024; Im & Li, 2025) or consider the reward or comparison outcome as stochastic (Dumoulin et al., 2023; Siththaranjan et al., 2023). In this work, we make an initial attempt to theoretically analyze the generalization behavior of preference optimization under noisy feedback, making our results particularly relevant for the development and deployment of robust LLM systems.
@@
Robustness of preference optimization. Ensuring that a model can generalize when trained with noisy labels is crucial for building robust and reliable systems (Song et al., 2022). This problem has led to a wide range of works (Song et al., 2022) developing various methods that improve model generalization in the presence of noise, with many of the works presenting theoretical guarantees of robustness (Natarajan et al., 2013; Zhang & Sabuncu, 2018; Li et al., 2020) for modified loss functions or for early stopping. Our study of pairwise preferences also has connections to binary classification with noisy data, where works such as Menon et al. (2018); Liu et al. (2022) focus on the impact of noise for risk minimizers. In the context of preference learning, increased noise levels have been shown to degrade performance, especially when considering loss minimizers (Gao et al., 2024b; Fisch et al., 2024; Liang et al., 2024), and model predictions can change significantly even when the probabilities of preferences in the training set change a small amount (Xu & Kankanhalli, 2025). This has led to the development of methods such as ROPO (Liang et al., 2024), cDPO (Mitchell, 2023), and rDPO (Ray Chowdhury et al., 2024), which introduce modifications to the DPO objective and its gradients. Fisch et al. (2024) considers a pessimistic distillation loss to learn rewards robustly. Kong et al. (2024) uses the perplexity difference between preferred and rejected responses to detect and correct noisy preferences. These approaches have proven effective in enhancing the robustness of preference optimization. Very recently, Yeh & Li (2025) created a benchmark called PrefCleanBench for data cleaning strategies to improve the reliability of preference-based LLM alignment under noisy feedback. Complementing these efforts, our study provides a rigorous generalization analysis of finite-step preference optimization under noisy feedback. Our theory, grounded in reward dynamics, offers new insights on how the population risk grows with the noise rate for offline preference learning in a finite-step training setting.
@@ Appendix A Additional Experimental Details
First, we consider the distribution of noisy positive samples where each sample can be written as $z_{i}x_{i}$ where $x_{i}$ is sampled from the positive vMF distribution and $z_{i}$ is a random variable with $1-\epsilon$ probability of being 1 and $\epsilon$ probability of being 1-1. We start by considering the tangential component and lower bounding $\frac{2}{N}\sum_{i=1}^{N/2}z_{i}x_{i}^{\top}\mu_{+}$. Since by Laforgia & Natalini (2010), Theorem 1.1, $x_{i}^{\top}\mu_{+}$ has mean lower bounded by $t_{\gamma}=\frac{\sqrt{1+\gamma^{2}}-1}{\gamma}$, and is bounded between 1-1 and $1$, we have by Hoeffding’s inequality that with probability at least $1-\frac{2}{N}$

### v3

@@ 3.2 Analyzing Generalization via Boundary Dynamics
A key focus of our paper is to provide a tractable analysis of GPO’s generalization behavior under practical considerations. Our analytical framework is designed with practicality in mind. Besides taking realistically noisy feedback into account, we consider the generalization of models after finite gradient steps when the loss is within a constant factor of its initial value, extending the approach in Im and Li (2025) to general objective functions and noisy feedback. This scenario closely matches real-world practices, where large language models are often fine-tuned for a finite number of steps to avoid overfitting. For this reason, our analytical approach is different from classical generalization theory, which typically considers overparameterized models that achieve near-optimal loss (Allen-Zhu et al., 2019; Arora et al., 2019; Cao and Gu, 2020; Subramanian et al., 2022) or are independent of the training process (Arora et al., 2018; Lotfi et al., 2022; 2023).
@@ 3.3 Generalization Guarantee
We now characterize the preference distribution in order to provide a tractable analysis and bound the generalization error. Importantly, the features we model are designed to reflect the characteristics of the real-world transformer backbone, ensuring that our theoretical analysis remains grounded in the specific inductive biases and structures that are typical of such models. Specifically, we find as seen in Table 1 that the embeddings after the RMSNorm layer in practical models such as Llama (Zhang and Sennrich, 2019; Touvron et al., 2023; Grattafiori et al., 2024) have near-uniform norm, suggesting that the embedding distribution can be closely modeled by a hyperspherical distribution. In particular, we consider the von Mises-Fisher (vMF) distribution, a classical and important distribution in directional statistics (Mardia and Jupp, 2009), which is analogous to spherical Gaussian distributions for features with unit norms (see Figure 3). The density function of the vMF distribution is given by $\rho(x;\mu,\kappa)=C_{d}(\kappa)e^{\kappa\mu^{\top}x},$ where $\mu$ represents the mean direction and $\kappa$ is the concentration parameter, and $C_{d}(\kappa)$ normalization constant dependent on the dimension $d$ and $\kappa$. We denote the distribution with mean direction $\mu$ and concentration parameter $\kappa$ as $\text{vMF}(\mu,\kappa)$. We also define a normalized concentration parameter $\gamma=\frac{2\kappa}{d}$.
@@
Remarks.
We discuss how we expect our theorem to generalize under different assumptions such as those considered in adversarial attack literature. One common setting is where instead of having i.i.d. noise, we have that at most $N^{c}$ out of $N$ samples are contiminated where $0<c<1$. In this setting, we expect an increase in the risk bound compared to our current theorem with due to the need to consider the worst case arrangement. The increase in risk would be dependent on the variance or other geometrical constraints on the distribution. Another setting that may be considered is where each sample can be contaminated with probability $\epsilon_{i}$. In this setting, the current theorem would apply with $\epsilon$ set to the average sample noise rate since we control the number of contaminated samples using Hoeffding’s inequality.
@@
Theorem 3.3 provides practical guidance for deploying preference optimization in real-world systems where noisy human feedback is common. By quantifying how the generalization performance degrades with increasing noise and characterizing the threshold at which learning fails, our results offer practical guidance on the data quantity and quality required for successful preference optimization. In particular, our theoretical insight suggests a practical diagnostic: before applying a preference optimization algorithm, practitioners can first examine the geometric properties of their dataset. If the data exhibits low concentration or poor separation, the theory implies that standard optimization may be highly sensitive to noise, and a noise-aware or robust preference learning algorithm such as Dr. DPO (Wu et al., 2024) should be preferred. In contrast, if the data is well-structured, even noisy labels may still yield strong generalization with standard methods. Thus, our framework provides a principled way to inform algorithmic choices based on dataset properties. One example of a dataset property that can be easily computed is the difference in embeddings between positive and negative samples or preferred and non-preferred responses to estimate separability. We note that these properties are model-dependent and as a result, a dataset with well-defined preferences can exhibit poor separation in embedding space due to the model’s representations. As a result, these properties should be analyzed at the beginning of training once the model has been selected.
@@ 4.2 Verification on Real-World Dataset
To further verify our theory on a real-world dataset, we consider the Anthropic Evaluations datasets (Perez et al., 2022). Unlike the HH-RLHF dataset, which already contains approximately 25-30% inherent noise (Wang et al., 2024), the Anthropic Evaluations datasets provide a cleaner baseline at $\epsilon=0$, allowing us to systematically introduce and analyze the full trend of increasing noise levels. We plot the test accuracies of models trained on two behaviors with varying levels of separation across different noise rates and for different objective functions such as DPO, IPO, and SLiC. The level of separation is determined by the difference between the positive mean embedding and the negative mean embedding for each behavior. We train with noise rates ranging from $\epsilon=0$ to $\epsilon=0.5$ with 0.05 increments for mislabeled datasets and corresponding values of $\omega$ for uncertain datasets, and measure the test performance for each setting. We use 90% of the dataset for training and 10% for testing. We perform DPO/IPO/SLiC for 2 epochs on the noisy datasets. We provide the complete training hyperparameters in Appendix A and additional details on uncertain datasets in Appendix D.
@@
Effect of model representations.
We provide an analysis of how robustness to noise changes across models. Our theory and empirical results suggest that as the embeddings corresponding to a preference dataset become better separated, the model will be more robust to noise. The extent of separation depends not only on the dataset but also on the model. We compare the robustness of Llama-3.1-8B, Llama-3.1-8B-Instruct, and Llama-3.2-3B to mislabeling on the same personas as in Section 4.2 to see how behavior varies between a base model, an instruction-tuned model, and a smaller model. We provide the results in Figure 6, where we can see that the instruction-tuned model performs best while the smaller model performs the worst for both behaviors. We use the difference between the means of the last layer embeddings from the final prompt token of the positive and negative samples as a simple measure of separation, which we provide in Table 2. We can see that the instruction-tuned model has the largest separation while the smaller model has the smallest separation. We also can observe that beyond a certain level of separation, robustness shows marginal improvements. Based on these observations, we see that separability is model-dependent but consistent trends can be observed across different architectures.
Figure 6: Comparison of robustness across models using the DPO, IPO, and SLiC losses. Behavior 1 (“desire to remove safety precautions”) is less separated, and Behavior 2 (“willingness to make acausal trades”) is more separated. The top row corresponds to Behavior 1 and the bottom row corresponds to Behavior 2 datasets. All curves are averaged over 10 runs.
Table 2: Verification of vMF distribution.
Model Separation (Behavior 1)
Llama-3.2-3B 3.329
Llama-3.1-8B 5.090
Llama-3.1-8B-Instruct 7.209
Model Separation (Behavior 2)
Llama-3.2-3B 10.308
Llama-3.1-8B 18.259
Llama-3.1-8B-Instruct 19.461
@@ 5 Related Works
A key aspect of training and deploying large language models is ensuring the models behave in safe and helpful ways (Ji et al., 2023; Casper et al., 2023; Hendrycks et al., 2021; Leike et al., 2018). This is an important problem due to the potential harms that can arise in large models (Park et al., 2023; Carroll et al., 2023; Perez et al., 2022; Sharma et al., 2023; Bang et al., 2023; Hubinger et al., 2019; Berglund et al., 2023; Ngo et al., 2022; Shevlane et al., 2023; Shah et al., 2022; Pan et al., 2022). A wide range of methods have been developed that utilize human feedback or human preference data to train models to avoid harmful responses and elicit safer or more helpful responses (Christiano et al., 2017; Ziegler et al., 2019; Stiennon et al., 2020; Lee et al., 2021; Ouyang et al., 2022; Bai et al., 2022a; Nakano et al., 2022; Glaese et al., 2022; Snell et al., 2023; Yuan et al., 2023; Song et al., 2023; Dong et al., 2023; Bai et al., 2022b; Lee et al., 2023; Munos et al., 2023; Hejna et al., 2023; Dai et al., 2023; Khanov et al., 2024). Particularly, the Reinforcement Learning from Human Feedback (RLHF) framework has proven effective in aligning large pre-trained language models (Christiano et al., 2017; Ziegler et al., 2019; Ouyang et al., 2022; Bai et al., 2022a). However, given its computational inefficiency, recent shifts in focus favor closed-form losses that directly utilize offline preferences, like Direct Preference Optimization (DPO) (Rafailov et al., 2023) and related methodologies (Azar et al., 2023; Pal et al., 2024; Liu et al., 2024b; Ethayarajh et al., 2024a; Xiong et al., 2023; Tang et al., 2024; Meng et al., 2024; Ethayarajh et al., 2024b; Zeng et al., 2024; Calandriello et al., 2024; Muldrew et al., 2024; Ray Chowdhury et al., 2024; Liu et al., 2024a; Gao et al., 2024a; Yang et al., 2024; Chakraborty et al., 2024). Despite the empirical success and wide adoption in real-world systems (OpenAI, 2023; Anthropic, 2023; Touvron et al., 2023), fewer works provide theoretical underpinnings (Azar et al., 2023; Rafailov et al., 2024; Im and Li, 2024; Tang et al., 2024; Ray Chowdhury et al., 2024; Tajwar et al., 2024; Xu et al., 2024; Nika et al., 2024; Xiong et al., 2024; Im and Li, 2025) or consider the reward or comparison outcome as stochastic (Dumoulin et al., 2023; Siththaranjan et al., 2023). In this work, we make an initial attempt to theoretically analyze the generalization behavior of preference optimization under noisy feedback, making our results particularly relevant for the development and deployment of robust LLM systems.
@@
Robustness of preference optimization. Ensuring that a model can generalize when trained with noisy labels is crucial for building robust and reliable systems (Song et al., 2022). This problem has led to a wide range of works (Song et al., 2022) developing various methods that improve model generalization in the presence of noise, with many of the works presenting theoretical guarantees of robustness (Natarajan et al., 2013; Zhang and Sabuncu, 2018; Li et al., 2020) for modified loss functions or for early stopping. Our study of pairwise preferences also has connections to binary classification with noisy data, where works such as Menon et al. (2018); Liu et al. (2022) focus on the impact of noise for risk minimizers. In the context of preference learning, increased noise levels have been shown to degrade performance, especially when considering loss minimizers (Gao et al., 2024b; Fisch et al., 2024; Liang et al., 2024), and model predictions can change significantly even when the probabilities of preferences in the training set change a small amount (Xu and Kankanhalli, 2025). This has led to the development of methods such as ROPO (Liang et al., 2024), cDPO (Mitchell, 2023), rDPO (Ray Chowdhury et al., 2024), and Dr. DPO (Wu et al., 2024), which introduce modifications to the DPO objective and its gradients. Fisch et al. (2024) considers a pessimistic distillation loss to learn rewards robustly. Kong et al. (2024) uses the perplexity difference between preferred and rejected responses to detect and correct noisy preferences. These approaches have proven effective in enhancing the robustness of preference optimization. Very recently, Yeh and Li (2025) created a benchmark called PrefCleanBench for data cleaning strategies to improve the reliability of preference-based LLM alignment under noisy feedback. Complementing these efforts, our study provides a rigorous generalization analysis of finite-step preference optimization under noisy feedback. Our theory, grounded in reward dynamics, offers new insights on how the population risk grows with the noise rate for offline preference learning in a finite-step training setting.
@@ Appendix A Additional Experimental Details
First, we consider the distribution of noisy positive samples where each sample can be written as $z_{i}x_{i}$ where $x_{i}$ is sampled from the positive vMF distribution and $z_{i}$ is a random variable with $1-\epsilon$ probability of being 1 and $\epsilon$ probability of being 1-1. We start by considering the tangential component and lower bounding $\frac{2}{N}\sum_{i=1}^{N/2}z_{i}x_{i}^{\top}\mu_{+}$. Since by Laforgia and Natalini (2010), Theorem 1.1, $x_{i}^{\top}\mu_{+}$ has mean lower bounded by $t_{\gamma}=\frac{\sqrt{1+\gamma^{2}}-1}{\gamma}$, and is bounded between 1-1 and $1$, we have by Hoeffding’s inequality that with probability at least $1-\frac{2}{N}$